# Configuring Agentic AI Coding Tools: An Exploratory Study

### 20 Mar 2026 v2

@@ Abstract.
journalyear: 2026copyright: ccdoi: XXXXXXX.XXXXXXXdoi: XXXXXXX.XXXXXXXconference: Anonymous Conference; Anoym. Conf.; Earthbooktitle: Anonymous Conferencedoi: 10.1145/XXXXXXX.XXXXXXXisbn: 979-8-4007-XXXX-X/2026/04ccs: Software and its engineering
@@ 1. Introduction
We define a configuration mechanism as a means for developers to tailor tool and agent behavior to a project or workflow (e.g., context files or dedicated subagents). A configuration artifact is a tangible instance of a mechanism (e.g., the actual Markdown file). For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2026; OpenAI, 2025). Configuration artifacts are version-controlled, inspectable, and collaboratively maintained. With the increased availability and diversity of agentic tools, the number of configuration mechanisms and related artifacts has also increased.
@@ 2. Related Work
Prior work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows, but has typically focused on single artifact types in isolation. There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior. Independent of tool architecture, context engineering allows developers to design, structure, and supply task- and project-specific information to guide the behavior of tools (Mei et al., 2025; Schmid, 2025; Horthy, 2025; Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025). Empirical work has mainly focused on repository-maintained context files, such as AGENTS.md. Mohsenimofidi et al. (Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level context files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions. Chatlatanagulchai et al. (Chatlatanagulchai et al., 2025) found that context files are actively maintained through frequent, small changes and focus predominantly on development instructions such as build and testing procedures. Santos et al. (Santos et al., 2025) analyzed 328 CLAUDE.md files and found that architectural concerns dominate, often co-occurring with development guidelines, testing, and project overview.
@@
Beyond structural characterization, recent work has begun to evaluate the impact of such artifacts. Lulla et al. (Lulla et al., 2026) conducted a controlled study comparing agent executions with and without an AGENTS.md file and reported lower runtime and token consumption while maintaining comparable task completion behavior.
@@
Villamizar et al. (Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda for their evolution, reuse, and governance. Zhang et al. (Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses brevity bias and context collapse via incremental updates.
@@
Recent work further generalized this perspective by formalizing context as a programmable abstraction. Zhang and Wang (Zhang and Wang, 2025) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition. McMillan (McMillan, 2026) evaluated structured file-native context schemas and showed how format and multi-file organization influence retrieval accuracy and downstream agent performance. Ye et al. (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search. Jiang and Nam (Jiang and Nam, 2026) explored Cursor rules as one particular configuration mechanism and developed a taxonomy of project contexts.
@@
However, a cross-tool perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
@@ 3. Data Collection and Analysis
We collected OSS projects hosted on GitHub, selecting repositories belonging to “engineered” software projects (Munaiah et al., 2017) using an updated version of a selection approach from prior work, starting with the SEART GitHub search tool (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021). To ensure sufficient project maturity and development history, we selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025. We then excluded archived, disabled, or locked repositories. We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026). This resulted in a sample of 37,249 repositories.
@@
For the remaining 36,184 repositories, we used a GPT-5.2-based classification pipeline to determine, based on README content, whether each project qualifies as “engineered” per Munaiah et al. (Munaiah et al., 2017), that is, showing clear evidence of software engineering practices (e.g., a stated purpose, installation and deployment instructions, documentation, testing, CI/CD) and maintenance. The supplementary material includes the prompt (iteratively developed and tested on subsets) and model configuration (OpenAI defaults) (Authors, 2026). The pipeline labeled 32,564 repositories as engineered and excluded the remaining 3,620 (1,416 classified as non-engineered, 2,204 “unsure”). We spot-checked randomly selected samples from each category during prompt development and for the final sample. We then cloned the remaining 32,564 repositories and applied heuristics based on file names and file paths to detect the usage of AI coding tools and configuration mechanisms (see Table 1). These heuristics are based on the answers to RQ1 and are briefly discussed in Section 4. This resulted in 2,923 repositories that use one or more AI coding tools (three additional repositories contained only empty configuration artifacts and were excluded). Our data collection and analysis scripts and the analyzed data are available online (Authors, 2026).
@@ 4. Configuration Mechanisms (RQ1)
We selected five agentic AI coding tools (Table 2) based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026): the four most popular AI tools among the surveyed developers, substituting Codex for ChatGPT as the agentic tool by the same vendor, plus Cursor, which has recently been studied in software engineering (He et al., 2026; Jiang and Nam, 2026). Our study focuses on the CLI-based agentic interfaces that first appeared in 2025; for some tools, prior non-agentic versions exist. One author systematically reviewed each tool’s online documentation, documenting configuration mechanisms together with repository-level files and directories that indicate their usage. This was then cross-checked by two other authors and discussed among all three.
We used these findings to develop heuristics for detecting tool usage and configuration mechanisms in repositories (Table 1). For example, a .claude directory or a CLAUDE.md file indicates Claude Code usage; a .claude/agents/ directory with Markdown files indicates Subagents. AGENTS.md is a special case: as a tool-agnostic standard, it is supported by multiple tools, so repositories using only AGENTS.md without tool-specific artifacts are tracked separately (see Table 2, footnote d). Note that for Copilot, Cursor, and Gemini, certain detected files apply to both conversational and agentic interfaces. All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Authors, 2026).
@@ 5.2. Distribution of Configuration Mechanisms
Figure 3 shows the distribution of configuration mechanisms per tool. Context Files (e.g., CLAUDE.md, AGENTS.md) were the most frequently adopted mechanism, used by 61.5–100% of repositories across all tools. Beyond Context Files, distinct tool-specific profiles emerge: 72.8% of Cursor repositories adopt Rules, since Cursor was one of the first to introduce this mechanism (Jiang and Nam, 2026), and 61.2% of Gemini repositories use Settings. Other mechanisms remain below 20% of repositories for most tools.
@@ 6.1. Configuration Mechanism: Context Files
Context Files are Markdown files that provide central, machine-readable context about a project. AGENTS.md, introduced by OpenAI, now serves as an open tool-agnostic convention increasingly supported by various tools (Mohsenimofidi et al., 2026). We identified 4,854 Context Files across 2,631 of the 2,923 repositories in our sample. Most repositories include one or two such files (Figure 5).
@@ 6.4. Threats to Validity
We classified repositories as “engineered” software projects using GPT-5.2 on their README content, with iterative prompt development and spot-checks to mitigate errors. We used a single labeling run without considering inter-model agreement; 2,204 “unsure” cases (often due to inaccessible linked resources) were excluded. Future work should assess the reliability of this approach by adding alternative models and configurations. Detection heuristics were designed by one author and cross-checked by two others, but rapidly evolving tool ecosystems may introduce conventions not yet captured.
@@
Our study covers only open-source repositories on GitHub; practices in proprietary or enterprise settings may differ. We selected repositories showing established engineering practices, but cannot claim representativeness for closed-source development, nor did we examine variation across application domains. Finally, our findings are a point-in-time snapshot (February 2026) of a rapidly evolving landscape.
@@ 7. Discussion
Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,923 repositories. Below, we discuss the implications of our study.
@@
Configuration practices vary systematically across tools: Claude Code repositories exhibit the broadest configuration footprint, Cursor projects emphasize Rules and Commands, and Copilot repositories rarely extend beyond Context Files. These differences likely reflect both the configuration options each tool exposes (Table 1) and the norms emerging within their respective user communities. Repository characteristics also differ: Cursor repositories tend to be younger and larger, while Gemini repositories show higher activity levels (Table 3). The 528 repositories using AGENTS.md without any tool-specific configuration further reinforce its role as a tool-agnostic standard. It remains an open question whether these tool-specific configuration cultures will converge as tool capabilities increasingly overlap, or whether distinct usage patterns will persist.
@@
For researchers, our findings point to several directions. Longitudinal studies are needed to track how configuration practices evolve as tools mature. Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone (Lulla et al., 2026). Given that certain projects use multiple tools (Section 5.1), research should also investigate interoperability challenges, including how conflicts between overlapping or contradictory instructions across files can be detected and resolved. Our automated data collection and analysis pipeline (Authors, 2026) supports such ongoing analyses.
Finally, our findings represent a point-in-time snapshot of a rapidly evolving landscape. The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.
@@ 8. Conclusions
Our study is the first to provide a comprehensive overview of the configuration mechanisms used by agentic AI coding tools in software engineering. We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,923 GitHub repositories. Three findings stand out. First, Context Files dominate and are often the only mechanism present, with AGENTS.md emerging as an interoperable standard. Second, advanced mechanisms such as Skills and Subagents remain shallowly adopted, with Skills used predominantly as static documents rather than executable workflows. Third, distinct configuration cultures are forming around different tools. For practitioners, these findings suggest that AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and that Skills offer untapped potential for encoding recurring workflows beyond static instructions. Tool providers should consider improving onboarding and documentation for advanced mechanisms, given the gap between their expressive power and current adoption. For researchers, controlled studies should determine whether richer configuration yields measurable improvements over Context Files alone, and longitudinal research should track how configuration patterns evolve.

### 09 Apr 2026 v3

@@ Abstract.
journalyear: 2026copyright: ccdoi: 10.1145/XXXXXXX.XXXXXXXconference: 3rd ACM International Conference on AI-powered Software; July 6–7, 2026; Montreal, Canadabooktitle: 3rd ACM International Conference on AI-powered Software (AIware 2026), July 6–7, 2026, Montreal, Canadaisbn: 979-8-4007-XXXX-X/2026/04ccs: Software and its engineering
@@ 1. Introduction
We define a configuration mechanism as a means for developers to tailor tool and agent behavior to a project or workflow (e.g., context files or dedicated subagents). A configuration artifact is a tangible instance of a mechanism (e.g., the actual Markdown file). For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2025; OpenAI, 2025). Configuration artifacts are version-controlled, inspectable, and collaboratively maintained. With the increased availability and diversity of agentic tools, the number of configuration mechanisms and related artifacts has also increased.
@@ 2. Related Work
Prior work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows, but has typically focused on single artifact types in isolation. There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior. Context engineering builds on earlier prompt engineering techniques (Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025) but broadens the scope from individual prompts to the systematic design of all information supplied to a tool. Mei et al. (Mei et al., 2025) survey this emerging field, while practitioner accounts demonstrate how structuring project-specific context improves agent effectiveness in complex codebases (Schmid, 2025; Horthy, 2025). Empirical work has mainly focused on repository-maintained context files, such as AGENTS.md. Mohsenimofidi et al. (Mohsenimofidi et al., 2025) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level context files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions. Chatlatanagulchai et al. (Chatlatanagulchai et al., 2025) found that context files are actively maintained through frequent, small changes and focus predominantly on development instructions such as build and testing procedures. Santos et al. (Santos et al., 2025) analyzed 328 CLAUDE.md files and found that architectural concerns dominate, often co-occurring with development guidelines, testing, and project overview.
@@
Beyond structural characterization, Lulla et al. (Lulla et al., 2026) conducted a controlled study comparing agent executions with and without an AGENTS.md file and reported lower runtime and token consumption while maintaining comparable task completion behavior.
@@
Villamizar et al. (Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda for their evolution, reuse, and governance. Q. Zhang et al. (Zhang et al., 2025a) proposed a framework treating contexts as evolving instructions, addressing brevity bias and context collapse through incremental updates.
@@
Other work formalizes context as a programmable abstraction. Y. Zhang et al. (Zhang et al., 2025b) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition. McMillan (McMillan, 2026) evaluated structured file-native context schemas and showed how format and multi-file organization influence retrieval accuracy and downstream agent performance. Ye et al. (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search. Jiang and Nam (Jiang and Nam, 2025) explored Cursor rules and developed a taxonomy of project contexts.
@@
However, no prior work maps configuration mechanisms across tools or studies their adoption patterns across repositories. We provide this cross-tool perspective.
@@ 3. Data Collection and Analysis
We collected OSS projects hosted on GitHub, selecting repositories belonging to “engineered” software projects (Munaiah et al., 2017) using an updated version of a selection approach from prior work, starting with the SEART GitHub search tool (Mohsenimofidi et al., 2025; SEART, 2025; Dabic et al., 2021). To ensure sufficient project maturity and development history, we selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025. We then excluded archived, disabled, or locked repositories. We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and excluded repositories with fewer than 271 commits or fewer than 7 watchers (median-based thresholds from prior research (Mohsenimofidi et al., 2025)). This resulted in a sample of 37,249 repositories.
@@
For the remaining 36,184 repositories, we used a GPT-5.2-based classification pipeline to determine, based on README content, whether each project qualifies as “engineered” per Munaiah et al. (Munaiah et al., 2017), that is, showing clear evidence of software engineering practices (e.g., a stated purpose, installation and deployment instructions, documentation, testing, CI/CD) and maintenance. The supplementary material includes the prompt (iteratively developed and tested on subsets) and model configuration (OpenAI defaults) (Galster et al., 2026). The pipeline labeled 32,564 repositories as engineered and excluded the remaining 3,620 (1,416 classified as non-engineered, 2,204 “unsure”). We spot-checked randomly selected samples from each category during prompt development and for the final sample. We then cloned the remaining 32,564 repositories and applied heuristics based on file names and file paths to detect the usage of AI coding tools and configuration mechanisms (see Table 1). These heuristics are based on the answers to RQ1 and are briefly discussed in Section 4. This resulted in 2,923 repositories that use one or more AI coding tools (three additional repositories contained only empty configuration artifacts and were excluded). Our data collection and analysis scripts and the analyzed data are available online (Galster et al., 2026).
@@ 4. Configuration Mechanisms (RQ1)
We selected five agentic AI coding tools (Table 2) based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026): the four most popular AI tools among the surveyed developers, substituting Codex for ChatGPT as the agentic tool by the same vendor, plus Cursor, which has recently been studied in software engineering (He et al., 2025; Jiang and Nam, 2025). Our study focuses on the CLI-based agentic interfaces that first appeared in 2025; for some tools, prior non-agentic versions exist. One author systematically reviewed each tool’s online documentation, documenting configuration mechanisms together with repository-level files and directories that indicate their usage. This was then cross-checked by two other authors and discussed among all three.
We used these findings to develop heuristics for detecting tool usage and configuration mechanisms in repositories (Table 1). For example, a .claude directory or a CLAUDE.md file indicates Claude Code usage; a .claude/agents/ directory with Markdown files indicates Subagents. AGENTS.md is a special case: as a tool-agnostic standard, it is supported by multiple tools, so repositories using only AGENTS.md without tool-specific artifacts are tracked separately (see Table 2, footnote d). Note that for Copilot, Cursor, and Gemini, certain detected files apply to both conversational and agentic interfaces. All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Galster et al., 2026).
@@ 5.2. Distribution of Configuration Mechanisms
Figure 3 shows the distribution of configuration mechanisms per tool. Context Files (e.g., CLAUDE.md, AGENTS.md) were the most frequently adopted mechanism, used by 61.5–100% of repositories across all tools. Beyond Context Files, distinct tool-specific profiles emerge: 72.8% of Cursor repositories adopt Rules, since Cursor was one of the first to introduce this mechanism (Jiang and Nam, 2025), and 61.2% of Gemini repositories use Settings. Other mechanisms remain below 20% of repositories for most tools.
@@ 6.1. Configuration Mechanism: Context Files
Context Files are Markdown files that provide central, machine-readable context about a project. AGENTS.md, introduced by OpenAI, now serves as an open tool-agnostic convention increasingly supported by various tools (Mohsenimofidi et al., 2025). We identified 4,854 Context Files across 2,631 of the 2,923 repositories in our sample. Most repositories include one or two such files (Figure 5).
@@ 6.4. Threats to Validity
We classified repositories as “engineered” software projects based on their README content. During iterative prompt development, we tested GPT-5.2, GPT-5-nano, and GPT-5-mini, and selected GPT-5.2 because spot-checks showed it produced fewer “unsure” labels and fewer false positives and negatives. We used a single labeling run; 2,204 “unsure” cases (often due to inaccessible linked resources) were excluded. Future work should assess the reliability of this approach by adding alternative models and configurations. Detection heuristics were designed by one author and cross-checked by two others, but rapidly evolving tool ecosystems may introduce conventions not yet captured.
@@
Our study covers only open-source repositories on GitHub; practices in proprietary or enterprise settings may differ. We selected repositories showing established engineering practices, but cannot claim representativeness for closed-source development, nor did we examine variation across application domains. Finally, our findings are a point-in-time snapshot (February 2026) of a fast-moving field.
@@ 7. Discussion
Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,923 GitHub repositories. Below, we discuss the implications of our findings.
@@
Configuration practices vary systematically across tools: Claude Code repositories exhibit the broadest configuration footprint, Cursor projects emphasize Rules and Commands, and Copilot repositories rarely extend beyond Context Files. These differences likely reflect the configuration options each tool exposes (Table 1) and the norms emerging within their user communities. Repository characteristics also differ: Cursor repositories tend to be younger and larger, while Gemini repositories show higher activity levels (Table 3). The 528 repositories using AGENTS.md without any tool-specific configuration further reinforce its role as a tool-agnostic standard. It remains an open question whether these tool-specific configuration cultures will converge as tool capabilities increasingly overlap, or whether distinct usage patterns will persist.
@@
For researchers, our findings motivate longitudinal studies to track how configuration practices evolve as tools mature, and controlled studies to assess whether Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone (Lulla et al., 2026). Given that certain projects use multiple tools (Section 5.1), research should also investigate how conflicts between overlapping instructions across files can be detected and resolved. Our data collection and analysis pipeline (Galster et al., 2026) supports such ongoing analyses.
Finally, our findings represent a point-in-time snapshot of a fast-moving field. The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.
@@ 8. Conclusions
Our study provides the first comprehensive overview of configuration mechanisms for agentic AI coding tools, identifying eight mechanisms across five tools and analyzing their adoption in 2,923 GitHub repositories. Three findings stand out. First, Context Files dominate and are often the only mechanism present, with AGENTS.md emerging as an interoperable standard. Second, advanced mechanisms such as Skills and Subagents remain shallowly adopted, with Skills used predominantly as static documents rather than executable workflows. Third, distinct configuration cultures are forming around different tools. For practitioners, AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and Skills offer untapped potential for encoding recurring workflows beyond static instructions. Tool providers should improve onboarding for Skills and Subagents, which few repositories use despite their ability to encapsulate executable workflows and isolated agent contexts. For researchers, controlled studies should determine whether richer configuration improves outcomes over Context Files alone, and longitudinal research should track how these patterns evolve.